All 3 questions are related to each other. Please answer all of them..
Experiment Lab 6-Post For the two titrations in a and b, calculate the KAvalue of each weak acid from the one halfway, and three-quarters points. monoprotic acid HQ 0.100 M NaoH titrating the weak Point in Titration volume (in mL) of NaOH pH 5.00 1/4 Point 4.57 10.00 5.05 1/2 Point 15.00 3/4 Point b. 0.0500 M NaOH titrating the weak monoprotic acid HP Volume (in mL) of Point in Titration NaOH 7.500 1/4 Point 15.00 3.96 1/2 Point 22.50 3/4 Point
media%2F1bd%2F1bd2ae0c-1964-447e-8827-f4
Answer #1

Q1.

Note that Ka value will NOT change due to volume added in titration, so ti is the same for all

pH = pKa + log(A-/HA)

in the half equivalence point A- = HA so

5.05= pKa + log(1)

pKa = 5.05

Ka = 10^-pKA = 10^-5.05 = 0.0000089125= 8.9125*10^-6

note that Ka value will be the same for 1/4 and 3/4 points

b)

M = 0.05 M so:

pH = pKa + log(A-/HA)

again

3.5 = pKa + log(1)

pKa = 3.5

Ka = 10^-PKa1 = 10^-3.5 =3162*10^-4

Q2.

Ka value should not change as we measuer it in 1/4, 2/4 or 3/4 values

If there is any change, is due to the measurement

Q3.

Ka value difference --> Volume addition, pH measurement, claculations, etc..